\begin{thebibliography}{99}

\bibitem[AG]{AG} Allenby, R. B. J. T. and Gregorac, R. J. (1973). On locally extended residually finite groups. In Conference on Group Theory (pp. 9–17). Springer Berlin
    Heidelberg.

\bibitem[AL]{AL}  K.~Almeida and I.~Lima (2021), Subgroup separability of Artin groups, Journal of Algebra, 583, 25--37.

\bibitem[B]{B} Burns, R. G. (1971). On Finitely Generated Subgroups of Free Products. Journal of the Australian Mathematical Society, 12(3), 358–364.

\bibitem[Ca]{Ca} McCammond, J. (2017). The mysterious geometry of Artin groups. Winter Braids Lecture Notes, 4, 1–30.

\bibitem[C]{C} Charney, R. (2000). The Tits conjecture for locally reducible Artin groups. International Journal of Algebra and Computation, 10(06), 783–797.

\bibitem[CP]{CP} Crisp, J. and Paris, L. (2001). The solution to a conjecture of Tits on the subgroup generated by the squares of the generators of an Artin group. Inventiones Mathematicae, 145(1), 19–36.

\bibitem[DM]{DM} Dasbach, O. and Mangum, B. (2001). The automorphism group of a free group is not subgroup separable. AMS IP studies in advanced mathematics, 24, 23-28.


\bibitem[D]{D} Deligne, P. (1972). Les immeubles des groupes de tresses généralisés. Inventiones Mathematicae, 17(4), 273–302.

\bibitem[G]{G} Gordon, C. M. (2004). Artin groups, 3-manifolds and coherence. Bol. Soc. Mat. Mexicana, 10, 193-198.


\bibitem[H]{H} Hall, M. (1949). Coset representations in free groups. Transactions of the American Mathematical Society, 67(2), 421–432.


  \bibitem[JS]{JS} Jankiewicz, Kasia, and Kevin Schreve. Right‐angled Artin subgroups of Artin groups. Journal of the London Mathematical Society 106.2 (2022): 818-854.

\bibitem[L]{L} Van der Lek, H. (1983). The homotopy type of complex hyperplane complements. Diss. Radboud University Nijmegen.

\bibitem[LR]{LR} Long, D. D. and Reid, A. W. (2008). Subgroup separability and virtual retractions of groups. Topology, 47(3), 137–159.

\bibitem[M]{M} Mal’cev, A. I. (1983). On homomorphisms onto finite groups. American Mathematical Society Translations, Series 2, 119, 67 – 79.

 \bibitem[Mi]{Mi} Minasyan, A. (2019). Virtual Retraction Properties in Groups. International Mathematics Research Notices, 1-44.

\bibitem[MR]{MR} Metaftsis, V. and Raptis, E. (2008). On the profinite topology of right-angled Artin groups. Journal of Algebra, 320(3), 1174–1181.

\bibitem[MT]{MT} Moldavanskii, D. and Timofeeva, L. (1987). Finitely generated subgroups of a group which is defined by one relation and has a nontrivial center that are finitely separable. Izvestiya Vysshikh Uchebnykh Zavedenii Matematika, 12, 58-59.

\bibitem[NW]{NW} Niblo, G. A. and Wise, D. T. (2000). Subgroup separability, knot groups and graph manifolds. Proceedings of the American Mathematical Society, 129(3), 685–693.

\bibitem[P]{P} Paris, L. (2014). $K(\pi ,1)$ conjecture for Artin groups. Annales de La Faculté Des Sciences de Toulouse Mathématiques, 23(2), 361–415.


\bibitem[S]{S} Scott, P. (1978). Subgroups of Surface Groups are Almost Geometric. Journal of the London Mathematical Society, s2-17(3), 555–565.

\bibitem[S2]{S2} Scott, P. (1985). Correction to ‘Subgroups of Surface Groups are almost Geometric.’ Journal of the London Mathematical Society, s2-32(2), 217–220.

\bibitem[Ser]{Ser} J.–P. SERRE. Arithmetic Groups. Homological group theory (Proc. Sympos., Durham, 1977),
London Math. Soc. Lecture Note Ser., vol. 36 (Cambridge University Press, Cambridge-New
York, 1979), pp. 105–136. MR 564421 (82b:22021).

\bibitem[W]{W} Wilton, H. (2008). Hall’s Theorem for Limit Groups. Geometric and Functional Analysis, 18(1), 271–303.

\bibitem[Wi]{Wi} Wise, D. T. (2006). Subgroup separability of the figure 8 knot group. Topology, 45(3), 421–463.

    \bibitem[Wi2]{Wi2} Wise, D. (2013). The last incoherent Artin group. Proceedings of the American Mathematical Society, 141(1), 139-149.

    \bibitem[Wi3]{Wi3} Wise, D. T. (2019). "An Invitation to Coherent Groups". What's Next?: The Mathematical Legacy of William P. Thurston (AMS-205), edited by Dylan Thurston, Princeton: Princeton University Press, pp. 326-414. https://doi.org/10.1515/9780691185897-014


\end{thebibliography}

\documentclass{article}

\usepackage[T1]{fontenc}
\usepackage{lmodern}
\usepackage{geometry}

\usepackage{amsthm,amsfonts}
\usepackage[sumlimits]{amsmath}
\usepackage{amssymb}
\usepackage[all]{xy}
\usepackage{tikz}
\usepackage{graphicx}
\usepackage{url}
\usepackage{tikz-cd}
\usepackage{pgf}
\usepackage{ragged2e}
\usepackage{hyperref}
\usepackage{tikz}
\usetikzlibrary{arrows,decorations.pathmorphing,backgrounds,positioning,fit,petri,calc}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{booktabs}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{lipsum}

\newtheorem{lema}{Lemma}[section]
\newtheorem{teo}[lema]{Theorem}

\newtheorem{prop}[lema]{Proposition}
\theoremstyle{definition}

\newtheorem*{defi}{Definition}

\numberwithin{equation}{section}

\newcommand{\cf}{\emph{cf.} }
\newcommand{\tq}{\,|\,}

\newcommand{\ed}{\ar@{-}}

\renewcommand{\epsilon}{\varepsilon}

\numberwithin{equation}{section}

\title{Subgroup Separability of Artin groups II}

\author{Kisnney Emiliano de Almeida\\Departamento de Ciências Exatas\\ Universidade Estadual de Feira de Santana (UEFS)\\ Av. Transnordestina S/N\\ 44036-900\\  Feira de Santana - BA\\  Brazil\\
\texttt{kisnney@gmail.com}\\ \and Igor Lima\\Departamento de Matemática\\ Universidade de Brasília (UnB)\\70910-900\\ Brasília - DF\\Brazil\\  \texttt{igor.matematico@gmail.com}}

\begin{document}
\maketitle

\abstract{In this paper, we establish a new criterion for determining whether an Artin group is subgroup separable (LERF), building upon a criterion introduced in a previous work. Specifically, we prove an Artin group is LERF if and only if it does not contain certain induced subgraphs, thus providing a more direct generalization of the Metaftsis-Raptis criterion for RAAGs. As consequences, we prove an Artin group is ERF if and only if it is a free abelian group and we establish a link between subgroup separability and coherence for Artin groups. }

\section{Introduction}

We commence by recalling that the cosets of finite index normal subgroups of a group $G$ form a basis for the profinite topology of $G$. A group $G$ is called {\bfseries Residually Finite} if this topology is Hausdorff, or equivalently, if the trivial subgroup is closed in it. That is equivalent to the intersection of all its finite index subgroups to be trivial, among other possible definitions.

There are some natural stronger versions of this property, and we will deal with two of them. We say a group $G$ is {\bfseries (Locally) Extended  Residually Finite}, or just {\bfseries (L)ERF}, if every (finitely generated) subgroup of $G$ is closed in its profinite topology. A LERF group is also called {\bfseries subgroup separable}, which is equivalent to every finitely generated subgroup of $G$ to be equal to an intersection of finite index subgroups of $G$.

Subgroup separability was originally introduced by Hall \cite{H}, possessing significant implications for both group theory and low-dimensional topology. Nevertheless, it has been affirmatively or negatively established for only a limited number of group classes. This property enables the demonstration that specific immersions can elevate to embeddings within a finite cover (\cite{S}, \cite{S2}, \cite{W}, \cite{LR}, \cite{Wi}). Moreover, for finitely presented groups, subgroup separability guarantees the solvability of the generalized word problem. This means that there exists an algorithm capable of determining, for every element $g \in G$ and any finitely generated subgroup $H \leq G$, whether $g$ is a member of $H$ or not \cite{M}. Subgroup separability is also correlated with other group properties (\cf \cite{Mi}, for example).

Artin groups represent a diverse class of groups, defined by a finite simplicial labeled graph. They encompass a wide range of groups, including free groups, free abelian groups, classical Braid groups, and various other subclasses. Many conjectures surround Artin groups, with most of them lacking definitive answers. For instance, there are ongoing conjectures establishing torsion-freeness, solvable word problem, centerlessness (for irreducible not spherical Artin groups), among other unsolved questions. Further details can be found at \cite{Ca}. 

Metaftsis and Raptis \cite[Theorem 2]{MR} established a Right-Angled Artin group is subgroup separable if and only if it does not contain certain induced subgraphs. In a previous work   \cite{AL} we have combined the Metaftis-Raptis criterion and Jankiewicz-Schreve results on the Generalized Tits Conjecture \cite{JS} to establish a general criterion for subgroup separability of all Artin groups. However, this general criterion has a different, more constructive shape. We will give more details on Artin groups and previous results on its subgroup separability in Section \ref{preliminaries}.

In this paper, we give a new criterion for subgroup separability of Artin groups that directly generalizes the Metaftsis-Raptis criterion for RAAGs. Through this paper, all graphs are finite simplicial and with edges labeled by integers greater than 1. We recall a subgraph $\Theta$ of a graph $\Gamma$ is called {\bfseries full} or {\bfseries induced}  if every edge of $\Gamma$ which connects vertices of $\Theta$ is also inside $\Theta$.

\begin{defi}\label{defpoisgr}
  We say a graph $\Gamma$ is \emph{poisonous} if it contains a full subgraph $\gamma$ that satisfies one of the following conditions:
  \begin{enumerate}
  \item
  $\gamma$ is a square of edges labeled by 2;

  \item
  $\gamma$ is a non-closed path of length three with all edges labeled by 2;

  \item
  $\gamma$ is connected with exactly three vertices and at most one edge labeled by 2;

  \item
  $\gamma$ is  isomorphic to the graph
 $$ \begin{tikzpicture}
  \node[circle,draw](a) {};
\node[circle,draw] (b) [right=of a] {};
  \node[circle,draw] (c) [below=of b] {};
  \node[circle,draw] (d) [below=of a] {};

  \draw[-] (a) -- node[label=above:$2$]{} (b);
  \draw[-] (b) -- node[label=right:$2$]{} (c);
    \draw[-] (c) -- node[label=below:$2$]{} (d);
      \draw[-] (d) -- node[label=left:$2$]{} (a);
        \draw[-] (a) -- node[label=above:$m$]{} (c);
      \end{tikzpicture},$$

    with $m>2$.

  \item
  $\gamma$ is  isomorphic to the graph

   $$ \begin{tikzpicture}
  \node[circle,draw](a) {};
\node[circle,draw] (b) [right=of a] {};
  \node[circle,draw] (c) [below=of b] {};
  \node[circle,draw] (d) [below=of a] {};

  \draw[-] (a) -- node[label=above:$2$]{} (b);
  \draw[-] (b) -- node[label=right:$2$]{} (c);
    \draw[-] (c) -- node[label=below:$2$]{} (d);
      \draw[-] (d) -- node[label=left:$2$]{} (a);
        \draw[-] (a) -- node[label=above:$m$,pos=0.4]{} (c);
        \draw[-] (b) -- node[label=below:$n$,pos=0.6]{} (d);
      \end{tikzpicture},$$

      with $m>2$ and $n>2$.
    \end{enumerate}

    For clarity, we will say a graph is \emph{$i$-poisonous} if it contains a full subgraph $\gamma$ that satisfies the $i$-th condition above.
\end{defi}

Note that if $\Gamma_1$ is a poisonous full subgraph of $\Gamma_2$ then $\Gamma_2$ is poisonous. We present our main result below.

\begin{teo}\label{teopois}
  Let $A=A(\Gamma)$ be an Artin group. Then $A$ is subgroup separable if and only if $\Gamma$ is not poisonous.
\end{teo}

From this result, we obtain two applications. The first one characterizes the ERF Artin groups, which are actually the finitely generated free abelian groups.

\begin{teo}\label{teoartinerf}
 Let $A=A(\Gamma)$ be an Artin group. Then the following are equivalent:
 \begin{enumerate}
   \item
   $A$ is ERF;

   \item
   $\Gamma$ is complete with all edges labeled by 2;

   \item
   $A$ is free abelian.
 \end{enumerate}
 \end{teo}

For the second application, we recall a group $G$ is called {\bfseries coherent} if every finitely generated subgroup of $G$ is finitely presented. It is well known that finite groups, free groups and fundamental groups of $2$ (or $3$)-manifolds, for instance, are coherent and coherence carries over to subgroups and quotients.
The group $F_2 \times F_2$ is not coherent, where $F_2$ is the free group of rank two, which implies $SL_n(\mathbb{Z})$ is not coherent for $n\geq 4$; it is an old open problem suggested by J.P. Serre in \cite[Problem F14]{Ser}  if $SL_3(\mathbb{Z})$ is coherent. A good survey on the subject may be found at \cite{Wi3}. 

The full description of coherent Artin groups is already stablished, as we will explain in Section 4. By using Theorem \ref{teopois} we can relate the two properties.

\begin{teo}\label{teocoh}
  Every subgroup separable Artin group is coherent. A coherent Artin group $A=A(\Gamma)$ is subgroup separable if and only if every non-closed full path of $\Gamma$ with length greater than 1 has length 2 with both edges labeled by 2.
\end{teo}

This paper is organized as follows: in Section 2 we explain some basic definitions and previously known results on the subject; in Section 3 we prove the main result; and in Section 4 we establish the applications of the main result.

\section{Preliminaries}\label{preliminaries}

Let $\Gamma$ be a finite simplicial graph, with edges labeled by integers greater than 1. Then the {\bfseries Artin group with  $\Gamma$ as underlying graph}, denoted by $A(\Gamma)$, is given by a finite presentation, with generators corresponding to the vertices of $\Gamma$ and relations given by

  \[  \underbrace{abab\cdots}_{m \text{ factors}}=\underbrace{baba\cdots}_{m \text{ factors}} \]

    for each edge of $\Gamma$ that connects the vertices $a$ and $b$ labeled by $m$. In that case we will say  $a$ and $b$ are {\bfseries $m$-adjacent}.

Let $V(\Gamma)$ be the set of vertices of $\Gamma$ and $E(\Gamma)$ be its set of edges. A subgraph $\gamma$ of $\Gamma$ is called a {\bfseries full} or {\bfseries induced} subgraph (generated by $V(\gamma)$) if every two vertices of $\gamma$ that are adjacent in $\Gamma$ are also adjacent in $\gamma$. We will denote the full subgraph generated by a subset $V_1$ of $V(\Gamma)$ by $S(V_1)$. We will also say $\gamma$ is a {\bfseries full path} of $\Gamma$ if it is a (non-closed) path and a full subgraph of $\Gamma$.

A subset $V_0\subset V(\Gamma)$ generates a subgroup of $A(\Gamma)$ that is isomorphic to $A(\Gamma_0$), where $\Gamma_0$ is the full subgraph of $\Gamma$ generated by $V_0$ \cite{L} - such subgroups are called {\bfseries parabolic subgroups} of $A(\Gamma)$.

We will strongly use the behaviour of subgroup separability under taking subgroups.

\begin{teo}[\cite{S},\cite{S2}]\label{teosublerf}
Let $H$ be a subgroup of $G$. If $H$ is not LERF then $G$ is not LERF either. The converse holds if $[G:H]<\infty$.
\end{teo}

Although it is not true in general that subgroup separability is preserved by semidirect products, it is true if one of the groups is ERF. 

\begin{teo}[\cite{AG}, Th. 4]\label{teoerf}
  Let $N$ be an ERF group and $M$ be a LERF group. Then any semidirect product $N\rtimes M$ is LERF.
\end{teo}

We define a {\bfseries 2-cone} of a graph $\Gamma'$ (with apex $u$), denoted by $\Gamma=\Gamma'+\{u\}$, as the graph $\Gamma$ with $V(\Gamma)=V(\Gamma')\sqcup \{u\}$ and such that $u$ is 2-adjacent in $\Gamma$ to every vertex of $\Gamma'$. Let $\mathcal{S}$ be the class of graphs  obtained from graphs with one or two vertices by taking disjoint unions and 2-cones.  The result  below is the known characterization of subgroup separable Artin groups.

\begin{teo}\label{teogeral}(\cite{AL})
  Let $A=A(\Gamma)$ be an Artin group, with $\Gamma$ as its underlying graph. Then $A$ is subgroup separable if and only if $\Gamma\in \mathcal{S}$.
\end{teo}

If $\Gamma$ is a graph, we also define
$$Z(\Gamma):=\{u\in V(\Gamma)\tq \text{$u$ is 2-adjacent to every $v\in V(\Gamma)$ such that $v\neq u$}\}$$

and recall the lemma below, which is essentially a particular case of Theorem \ref{teogeral}.

\begin{lema}\label{lemacongamma}(\cite{AL}, Lemma 4.6)
 Let $A=A(\Gamma)$ be a LERF Artin group such that $\Gamma$ is connected and has at least three vertices. Then $Z(\Gamma)\neq \emptyset$.
\end{lema}

\section{Proof of Theorem \ref{teopois}}

 Suppose $\Gamma$ is poisonous. Then $\Gamma$ contains a full subgraph $\gamma$ that satisfies one of the conditions of Definition \ref{defpoisgr}. Then $A(\gamma)$ is not LERF by Lemma \ref{lemacongamma} hence $A$ is not LERF by Theorem \ref{teosublerf}.

  Now suppose $\Gamma$ is not poisonous and let $V=V(\Gamma)$. We will proceed by induction on $|V|$ to prove that $A$ is LERF. By Theorem \ref{teogeral} it is enough to prove $\Gamma \in \mathcal{S}$.

  We will use the following notation: if $X\subset V$ then $S(X)$ is the full subgraph of $\Gamma$ generated by $X$.

  If $|V|$ is 1 or 2 then $\Gamma\in \mathcal{S}$ by definition. If $|V|=3$, then either $\Gamma$ is disconnected or $\Gamma$ is a 2-cone - either way, $\Gamma\in \mathcal{S}$.

   Let $n>3$, suppose the result is true for $|V|<n$ and let $A(\Gamma)$ be an Artin group such that $|V|=n$.

   Note that every non-empty proper full subgraph of $\Gamma$ belongs to $\mathcal{S}$. That is true because if $\Gamma'$ is a non-empty proper full subgraph of $\Gamma$ since $\Gamma$ is not poisonous then so is $\Gamma'$, then $A(\Gamma')$ is LERF by induction hypothesis hence $\Gamma'\in \mathcal{S}$ by Theorem \ref{teogeral}. We will freely use that fact during the proof. That also means it is enough to prove  $Z(\Gamma)\neq \emptyset$ because if that's the case then $\Gamma$ decomposes as a 2-cone of a proper full subgraph, which belongs to $S$ by the argument above.

 Suppose $\Gamma$ is disconnected. Then $\Gamma=\Gamma_1\cup \Gamma_2$ such that $\Gamma_1$ and $\Gamma_2$ are non-empty proper full subgraphs of $\Gamma$. Then $\Gamma_1,\Gamma_2\in \mathcal{S}$ so $\Gamma\in \mathcal{S}$.

 So we may assume $\Gamma$ is connected.

  Let $v\in V$ and let $\Gamma_0=S(V\setminus \{v\})$. Note that $\Gamma_0\in \mathcal{S}$, since it is a non-empty proper full subgraph of $\Gamma$.

  Now suppose $\Gamma_0$ is disconnected, let's say $\Gamma_0=\Gamma_{01}\cup \cdots \cup \Gamma_{0k}$, with $\Gamma_{0i}$ connected for all $i$ and $k\geq 2$. Since $\Gamma$ is connected, then $v$ is adjacent to $\Gamma_{0i}$ for all $i$, let's say the edge $e_i$ connects $v$ to a vertex $u_{0i}$ of $\Gamma_{0i}$. Then the edges $e_i$ are all labeled by 2, otherwise $\Gamma$ would be $3$-poisonous. If there was a vertex of $\Gamma_{0i}$ which is $m$-adjacent to $v$, with $m>2$, then $\Gamma$ would also be $3$-poisonous, so every vertex of $\Gamma_0$ that is adjacent to $v$ is actually 2-adjacent to $v$. If all the vertices of $\Gamma_0$ are 2-adjacent to $v$ then $v\in Z(\Gamma)$ hence $\Gamma\in \mathcal{S}$. That means we may assume there is a vertex $w\in \Gamma_{0i}$ for some $i$ such that $w$ is not adjacent to $v$. Assume, without loss of generality, that $w\in \Gamma_{01}$. Note that $w\neq u_{01}$. 

  Suppose $w$ is adjacent to $u_{01}$. Then $w$ is actually 2-adjacent to $u_{01}$ otherwise $\Gamma$ would be 3-poisonous with $\gamma=S(w,u_{01},v)$. However that would mean $\Gamma$ is 2-poisonous, with $\gamma=S(w,u_{01},v,u_{02})$. That implies $w$ is not adjacent to $u_{01}$. Since $\Gamma_{01}$ is connected, that means $\Gamma_{01}$ has at least three vertices. On the other hand, $A(\Gamma_{01})$ is LERF since $\Gamma_{01}$  is a non-empty proper full subgraph of $\Gamma$. By Lemma \ref{lemacongamma} there is a vertex $u_{01}'\in Z(\Gamma_{01})$. Since $u_{01}'$ is 2-adjacent to $w$ which is not adjacent to $u_{01}$, we have $u_{01}\neq u_{01}' \neq w$. Then $u_{01}'$ is adjacent (hence 2-adjacent) to $v$, otherwise $\Gamma$ would be 2-poisonous with $\gamma=S(u_{01}',u_{01},v,u_{02})$. But that would also imply $\Gamma$ is 2-poisonous, with $\gamma=S(w,u_{01}',v,u_{02})$, which is a contradiction.

 So we may also assume $\Gamma_0$ is connected. Since $\Gamma_0$ is a proper full subgraph of $\Gamma$ with $|V|-1\geq 3$ vertices, by Lemma \ref{lemacongamma} there is a vertex $u\in Z(\Gamma_0)$, such that $\Gamma_0=\Gamma_1+\{u\}$. Let $V(\Gamma_1)=\{w_1,w_2,\dots, w_k\}$, with $k=|V|-2\geq 2$. 

If $u$ and $v$ are 2-adjacent, then $u\in Z(\Gamma)$ hence $\Gamma\in \mathcal{S}$. So we may assume $u$ and $v$ are not 2-adjacent.

Since $S(u,v,w_i)$ is not 3-poisonous for all $i$ then $w_i$ is 2-adjacent to $v$ for all $i$. Let $1<j\leq k$. Since $S(w_1,w_j,u,v)$ is not 1-poisonous, 4-poisonous nor 5-poisonous then $w_1$ is 2-adjacent to $w_j$. That implies $w_1\in Z(\Gamma)$ hence $\Gamma\in \mathcal{S}$.

\section{Applications}

\begin{proof}[Proof of Theorem \ref{teoartinerf}]
  If $\Gamma$ is complete with all edges labeled by 2 then $A$ is free abelian by definition. If $A$ is free abelian then $A$ is LERF; since $A$ is finitely generated abelian then every subgroup of $A$ is finitely generated hence $A$ is ERF.

  Suppose $A$ is ERF. Now suppose $\Gamma$ is not complete with all edges labeled by 2. Then there is a pair of vertices $u,v\in V=V(\Gamma)$ such that $u$ and $v$ are not 2-adjacent.

  Consider a new graph $\Gamma_1$ such that $V(\Gamma_1)=V\cup \{w_1,w_2\}$ and the edges of $\Gamma_1$ are:
  \begin{itemize}
    \item
    the edges of $\Gamma$;

    \item
    edges $\{\alpha, w_i\}$ labeled by 2 for each $\alpha\in V$ and $i=1,2$;

    \item
    an edge $\{w_1,w_2\}$ labeled by 3.
  \end{itemize}

  By the definition of an Artin group, it is easy to see that $A_1=A(\Gamma_1)\simeq A\times I$, where $I$ is the subgroup of $A_1$ generated by $w_1$ and $w_2$. Note that $I\simeq A(e)$, where $e$ is an edge labeled by 3.

  By Theorem \ref{teogeral}, $I$ is LERF hence $A_1$ is LERF by Theorem \ref{teoerf}. However $S(u,v,w_1,w_2)$ is either 4-poisonous (if $u$ and $v$ are not adjacent) or 5-poisonous (if $u$ and $v$ are $m$-adjacent for $m>2$), hence $\Gamma_1$ is poisonous. That contradicts Theorem \ref{teopois}.
  \end{proof}

We recall a graph is called \emph{chordal} if it contains no closed path of length greater than or equal to 4 as a full subgraph. The result below was rewritten under our notation.

\begin{teo}[\cite{G,Wi2}]\label{teoartincoh}
An Artin group $A=A(\Gamma)$ is coherent if and only if
$\Gamma$ is chordal, every complete subgraph of $\Gamma$ with 3 or 4 vertices has at most one edge label $>2$, and $\Gamma$ is not 4-poisonous.
\end{teo}

We will split the proof of Theorem \ref{teocoh} into the two next results.

\begin{prop}
Every subgroup separable Artin group is coherent.
\end{prop}

\begin{proof}
  Let $A=A(\Gamma)$ be a subgroup separable Artin group.  By Theorem \ref{teopois}, $\Gamma$ is not poisonous. Let $\gamma$ be a full subgraph of $\Gamma$.

  First we prove $\Gamma$ is chordal. Suppose $\gamma$ is a closed path of length $l$ greater or equal to 4. If all edges of $\gamma$ are labeled by 2 then either $\Gamma$ is 1-poisonous (if $l=4$) or 2-poisonous (if $l>4$), which is a contradiction. Then $\gamma$ contains an edge labeled by $m>2$, generated by vertices $u,v$. Since $\gamma$ is connected, then $v$ is adjacent to another vertex $w$ of $\gamma$. Since $l\geq 4$, then $S(u,v,w)$ is 3-poisonous and so is $\Gamma$, which is a contradiction. Then there can be no such $\gamma$ hence $\Gamma$ is chordal.

Now suppose $\gamma$ is complete.

If $\gamma$ has three vertices, since $\Gamma$ is not 3-poisonous then $\gamma$ has at most one edge label $>2$.

If $\gamma$ has four vertices, suppose it contains two edges labeled by $m>2$ and $n>2$. Those edges may not have a vertex in common, otherwise $\Gamma$ would be 3-poisonous. For the same reason, all other edges of $\gamma$ need to be labeled by 2. That means $\Gamma$ is 5-poisonous, which is a contradiction.

Since $\Gamma$ is obviously not 4-poisonous, then $A$ is coherent by Theorem \ref{teoartincoh}.
\end{proof}

\begin{prop}
Let $A=A(\Gamma)$ be a coherent Artin group. Then $A$ is subgroup separable if and only if every non-closed full path of $\Gamma$ which has length greater than 1  has length 2 with both edges labeled by 2.
\end{prop}

\begin{proof}
  Let $A=A(\Gamma)$ be a coherent Artin group.

   First suppose every non-closed full path of $\Gamma$ with length greater than 1  has length 2 with both edges labeled by 2. By Theorem \ref{teopois} it is enough to prove $\Gamma$ is not poisonous. The condition above guarantees that $\Gamma$ is not 2-poisonous. By Theorem \ref{teoartincoh}, $\Gamma$ is chordal hence not 1-poisonous. The last two conditions of the same theorem also tell us $\Gamma$ is not 4-poisonous nor 5-poisonous.

Suppose $\Gamma$ is 3-poisonous. Then there is a full subgraph $\gamma$ of $\Gamma$ which has exactly three vertices and at most one edge labeled by 2. If $\gamma$ is complete, that contradicts Theorem \ref{teoartincoh}; if it is not complete then that contradicts our hypothesis. Hence $\Gamma$ is not poisonous.

  Now suppose $A$ is subgroup separable. By Theorem \ref{teopois} $\Gamma$ is not poisonous. Let $\gamma$ be a non-closed full path of $\Gamma$ of length greater than 1. First suppose $\gamma$ contains an edge labeled by $m>2$ generated by vertices $u,v$. Since $\gamma$ is not an edge then there is a third vertex $w$ which is adjacent to $u$ or $v$. That means $S(u,v,w)$ is 3-poisonous and so is $\Gamma$, which is a contradiction. Then $\gamma$ has all edges labeled by 2. That implies the length of $\gamma$ is equal to 2, otherwise $\Gamma$ would be 2-poisonous, which concludes the proof.
\end{proof}

{\bfseries Acknowledgments}: We thank A. Krasilnikov and P. Zalesskii for the insightful questions that inspired this work. The first author was partially supported by FINAPESQ-UEFS, FAPDF, Brazil. The second author was partially supported by DPI/UnB, FAPDF, Brazil.

\end{document}